# A Comprehensive Study of Governance Issues in Decentralized Finance Applications

### 03 Jan 2024 v2

@@ Abstract.
Decentralized Finance (DeFi) is a prominent application of smart contracts, representing a novel financial paradigm in contrast to centralized finance. While DeFi applications are rapidly emerging on mainstream blockchain platforms, their quality varies greatly, presenting numerous challenges, particularly in terms of their governance mechanisms. In this paper, we present a comprehensive study of governance issues in DeFi applications. Drawing upon insights from industry reports and academic research articles, we develop a taxonomy to categorize these governance issues. We collect and build a dataset of 4,446 audit reports from 17 Web3 security companies, categorizing their governance issues according to our constructed taxonomy. We conducted a thorough analysis of governance issues and identified vulnerabilities in governance design and implementation, e.g., voting sybil attack and proposal front-running. Our findings highlight a significant observation: the disparity between smart contract code and DeFi whitepapers plays a central role in these governance issues. As an initial step to address the challenges of code-whitepaper consistency checks for DeFi applications, we built a machine-learning-based prototype, and validated its performance on eight widely used DeFi projects, achieving a 56.14% F1 score and a 80% recall. Since the development and evolution process of DeFi applications has some unique characteristics that are different from traditional software systems, namely decentralization, transparency, and changement via the governance mechanisms, we should considering how to build a trustworthy DeFi application by adapting the current software development paradigms or inventing a new approach. Our study culminates in providing several key practical implications for various DeFi stakeholders, including developers, users, researchers, and regulators, aiming to deepen the understanding of DeFi governance issues and contribute to the robust growth of DeFi systems.
@@ 1. Introduction
A crucial point to highlight from our study is that our findings offer several significant insights for multiple DeFi stakeholders. For researchers in software engineering, there is a need to study and develop governance frameworks and theories for DeFi, as well as methods for validating these governance systems. As an important application of blockchain, DeFi applications have a standardized governance system development process that is crucial. As a centralized software system, DeFi’s life cycle management is different from traditional software systems. In the traditional software life cycle, developers collect user needs and feedback and constantly make changes to the project. Changes in DeFi projects no longer follow this paradigm. The evolution of DeFi projects is defined by its governance system. All stakeholders in DeFi should have the power to determine how to evolve DeFi. Therefore, software engineers should solve problems in DeFi governance and propose a general framework for DeFi governance. For developers, it is essential to be cognizant of governance issues and to design transparent, ethical, and robust governance systems. For users and investors, they should investigate the governance structure of DeFi projects, such as the contract ownership, user privileges, and token power distribution. For regulators, it is crucial to both oversee governance structures and examine whitepapers of DeFi projects, as they are key to identify fradulent activities.
@@ 2. Methodology
Figure 1. Overview of our study methodology
@@ 2.1. Understanding DeFi Governance Concepts
We commenced our literature search with a systematic exploration of prominent databases, namely IEEE Xplore (iee, 2023), ACM Digital Library (acm, 2023), and Scopus (sco, 2023). We first defined the keywords that are directly related to the domain we were going to study, “blockchain governance”, “smart contract governance”, and “DeFi governance”. These predefined keywords were used to search the related articles. This initial phase yielded a substantial corpus of articles: 458 from IEEE Xplore, 498 from ACM Digital Library, and 1898 from Scopus, all published between 2017 and 2023 as shown in Figure 2. We can see that the focus on governance continues to increase. We should notice that the number for 2023 is not all. Recently, researchers have studied how GPT4 can be used to analyze documents (Mellon et al., 2022; Shrestha et al., 2023). To distill this vast collection, we also employed GPT4 as a query assistant,focusing specifically on the titles and abstracts to identify the 100 most pertinent articles from each database, with an emphasis on governance in decentralized finance applications. We upload the files that contain the title and abstracts to GPT4. Then, we instruct GPT4 to return the most related items with the prompt, ”according to the ’title’ and ’abstract’, please use the topic model to extract the top 100 items from the file that related to the topic ’Decentralization Finance Application Governance’ in a concise way.” Expanding our search horizon, we also utilized Google Scholar and Connected Papers (con, 2023), reviewing the first five pages of the search results to identify articles closely aligned with our research theme. We checked their titles and abstracts to see if they are related to decentralized finance and governance. In the end, this comprehensive approach led to the selection of 34 academic articles.
@@
In our quest for a holistic understanding of DeFi governance, we extended our exploration beyond academic literature to include industry perspectives. Recognizing the critical insights that Web3 entities offer, we expanded our dataset to encompass 11 blog articles about DeFi governance from leading firms (anonymous, 2023) in this field, like OpenZeppelin. This inclusion of diverse, industry-specific perspectives ensures a more comprehensive understanding of DeFi governance. Notably, during searching articles, we discovered that several esteemed international organizations have published reports on decentralization finance from EUROFI (Eurofi, 2022), WIFPR (WIFPR, 2021), OECD (OECD, [n. d.]), BIS (BIS, [n. d.]) and DBC (Erik Weitenberg, TNO). We use Google engine and the aforementioned keywords to search for these reports. We consider only the internal organization or the national institute. This led us to incorporate five such reports into our dataset, enriching our analysis with practical viewpoints. We used Google to collect industry blogs and organization reports, employing DeFi governance as a keyword search, and checked the first 10 pages. We have published the list of search results online (anonymous, 2023).
@@ 2.2. Analyzing DeFi Governance in Audit Reports
Figure 3. Severity and Status of the Issues in our dataset
@@ 3.1.1. Developing a Taxonomy for DeFi Governance
Figure 4. Three-Stage Development Process of DeFi Governance
@@ 3.2.1. Distribution of Governance Issues
Table 2. Raw Keywords for Issue Classification
Category Subcategory Keywords Citations
Governance Mechanism Governance Token governance token, vote, proposal, deceision-making, tally, abstation, quorum, veto (Dursun and Üstündağ, 2021; Beck et al., 2018; Ekal and Abdul-wahab, 2022; Sun et al., 2022; Liu et al., 2023a; Schneider et al., 2020; Baninemeh et al., 2023; Bhambhwani, 2022; Sun et al., 2022; Santana and Albareda, 2022; Jensen et al., 2021a)
Ownership owner, ownership, privilege
Tokenomics Utility Token supply, token distribution, token name, token usage, asset token, token utility (Pelt et al., 2021; Bhambhwani, 2022; Beck et al., 2018; Chaudhary et al., 2023; Jensen et al., 2021b; Kiayias and Lazos, 2022; Balcerzak et al., 2022; Kaal, 2020; Hunhevicz et al., 2022; Fiorentino and Bartolucci, 2021; Beck et al., 2018)
Revenue Stream transaction fee, trading fee, marketplace fee, borrow rate, protocol fee, premium fee, performance fee, token issuance, generic fee, interest rate, charge a fee
Incentive Mechanism lock up, total value locked, yield, borrow, airdrop, burn, stake, liquidity, lend, loan, referral, mint, incentive
Codebase code update contract, upgradable (Erik Weitenberg, TNO; ope, 2023; Ethereum, 2023; oce, 2019; governance, 2023; Polygon, 2023)
Figure 6. Overlapping of Different Categories about Governance Issues
@@ 3.2.2. Hierarchy of Governance Vulnerability
Figure 7. Hierarchy of Governance Issue
@@ 3.2.3. Common Topics on Governance Issues
Table 5. Top-3 Topic Words in all governance issues
@@
Gov Token proposal, contract, token voting power, power moved governance, macilious, lack event
Ownership centralization risk trust issue admin, keys blacklisted contract
Utility Token token distribution initial total supply, wrong total, supply restriction rewards, price, users, duplicate, liquidity
Incentive owner privileges, centralization risk mint token trust issue admin, keys
Revenue protocol fee, transaction interest rate, borrow, manipulate incompatibility deationary tokens
Codebase upgradeable contracts, storage centralized control contract, contract upgrade -
@@ 3.3.2. Detection of Inconsistency
Table 6. DeFi Tokenomics Configuration Inconsistency Checking, and F1 = 56.14%

### 11 Jan 2024 v3

@@ Abstract.
Decentralized Finance (DeFi) is a prominent application of smart contracts, representing a novel financial paradigm in contrast to centralized finance. While DeFi applications are rapidly emerging on mainstream blockchain platforms, their quality varies greatly, presenting numerous challenges, particularly in terms of their governance mechanisms. In this paper, we present a comprehensive study of governance issues in DeFi applications. Drawing upon insights from industry reports and academic research articles, we develop a taxonomy to categorize these governance issues. We collect and build a dataset of 4,446 audit reports from 17 Web3 security companies, categorizing their governance issues according to our constructed taxonomy. We conducted a thorough analysis of governance issues and identified vulnerabilities in governance design and implementation, e.g., voting sybil attack and proposal front-running. Our findings highlight a significant observation: the disparity between smart contract code and DeFi whitepapers plays a central role in these governance issues. As an initial step to address the challenges of code-whitepaper consistency checks for DeFi applications, we built a machine-learning-based prototype, and validated its performance on eight widely used DeFi projects, achieving a 56.14% F1 score and a 80% recall. The process of developing and evolving DeFi applications presents distinct traits that set it apart from conventional software systems. These include aspects like decentralization, transparency, and modification through governance mechanisms. Therefore, it’s essential to explore whether adapting existing software development paradigms or creating a new methodology is more effective in constructing trustworthy DeFi applications. Our study culminates in providing several key practical implications for various DeFi stakeholders, including developers, users, researchers, and regulators, aiming to deepen the understanding of DeFi governance issues and contribute to the robust growth of DeFi systems.
copyright: acmcopyrightjournalyear: 2018doi: XXXXXXX.XXXXXXX
@@ 1. Introduction
A crucial point to highlight from our study is that our findings offer several significant insights for multiple DeFi stakeholders. For researchers in software engineering, there is a need to study and develop governance frameworks and theories for DeFi, as well as methods for validating these governance systems. As an important application of blockchain, DeFi applications have a standardized governance system development process that is crucial. Unlike centralized software systems, the lifecycle management of DeFi diverges from traditional software approaches. In conventional software development, user needs and feedback drive continuous modifications. However, DeFi projects deviate from this model; their evolution is governed by their distinct governance systems. Stakeholders in DeFi should possess the authority to guide its evolution. Hence, it’s imperative for software engineers to address challenges within DeFi governance and contribute towards establishing a comprehensive framework for its governance. For developers, it is essential to be cognizant of governance issues and to design transparent, ethical, and robust governance systems. For users and investors, they should investigate the governance structure of DeFi projects, such as the contract ownership, user privileges, and token power distribution. For regulators, it is crucial to both oversee governance structures and examine whitepapers of DeFi projects, as they are key to identify fradulent activities.
@@ 2. Methodology
Figure 1. Overview of our study methodology.
@@ 2.1. Understanding DeFi Governance Concepts
We commenced our literature search with a systematic exploration of prominent databases, namely IEEE Xplore (iee, 2023), ACM Digital Library (acm, 2023), and Scopus (sco, 2023). We first defined the keywords that are directly related to the domain we were going to study, “blockchain governance”, “smart contract governance”, and “DeFi governance”. These predefined keywords were used to search the related articles. This initial phase yielded a substantial corpus of articles: 458 from IEEE Xplore, 498 from ACM Digital Library, and 1898 from Scopus, all published between 2017 and 2023 as shown in Figure 2. It is evident that there is a growing emphasis on governance. It is important to acknowledge that the data for 2023 does not represent the entire year. Recently, researchers have studied how GPT4 can be used to analyze documents (Mellon et al., 2022; Shrestha et al., 2023). To distill this vast collection, we also employed GPT4 as a query assistant,focusing specifically on the titles and abstracts to identify the 100 most pertinent articles from each database, with an emphasis on governance in decentralized finance applications. We upload the files that contain the title and abstracts to GPT4. Then, we instruct GPT4 to return the most related items with the prompt, ”according to the ’title’ and ’abstract’, please use the topic model to extract the top 100 items from the file that related to the topic ’Decentralization Finance Application Governance’ in a concise way.” Expanding our search horizon, we also utilized Google Scholar and Connected Papers (con, 2023), reviewing the first five pages of the search results to identify articles closely aligned with our research theme. We checked their titles and abstracts to see if they are related to decentralized finance and governance. In the end, this comprehensive approach led to the selection of 34 academic articles.
@@
In our quest for a holistic understanding of DeFi governance, we extended our exploration beyond academic literature to include industry perspectives. Recognizing the critical insights that Web3 entities offer, we expanded our dataset to encompass 11 blog articles about DeFi governance from leading firms (anonymous, 2023) in this field, like OpenZeppelin. This inclusion of diverse, industry-specific perspectives ensures a more comprehensive understanding of DeFi governance. Notably, during searching articles, we discovered that several esteemed international organizations have published reports on decentralization finance from EUROFI (Eurofi, 2022), WIFPR (WIFPR, 2021),OECD (OECD, [n. d.]),BIS (BIS, [n. d.]) and Dutch Blockchain Coalition (Erik Weitenberg, TNO).We use Google engine and the aforementioned keywords to search for these reports. We consider only the internal organization or the national institute. This led us to incorporate five such reports into our dataset, enriching our analysis with practical viewpoints. We used Google to collect industry blogs and organization reports, employing DeFi governance as a keyword search, and checked the first 10 pages. We have published the list of search results online (anonymous, 2023).
@@ 2.2. Analyzing DeFi Governance in Audit Reports
Figure 3. Severity and Status of the Issues in our dataset.
@@ 3.1.1. Developing a Taxonomy for DeFi Governance
Figure 4. Three-Stage Development Process of DeFi Governance.
@@ 3.2.1. Distribution of Governance Issues
Table 2. Raw Keywords for Issue Classification.
Category
Subcategory
Keywords
Citations
Governance Mechanism Governance Token
governance token, vote, proposal, deceision-making, tally, abstation, quorum, veto
(Dursun and Üstündağ, 2021; Beck et al., 2018; Ekal and Abdul-wahab, 2022; Sun et al., 2022; Liu et al., 2023a; Schneider et al., 2020) (Baninemeh et al., 2023; Bhambhwani, 2022; Sun et al., 2022; Santana and Albareda, 2022; Jensen et al., 2021a)
Ownership
owner, ownership, privilege
Tokenomics Utility Token
supply, token distribution, token name, token usage, asset token, token utility
(Pelt et al., 2021; Bhambhwani, 2022; Beck et al., 2018; Chaudhary et al., 2023; Jensen et al., 2021b; Kiayias and Lazos, 2022) (Balcerzak et al., 2022; Kaal, 2020; Hunhevicz et al., 2022; Fiorentino and Bartolucci, 2021; Beck et al., 2018)
Revenue Stream
transaction fee, trading fee, marketplace fee, borrow rate, protocol fee, premium fee, performance fee, token issuance, generic fee, interest rate, charge a fee
Incentive Mechanism
lock up, total value locked, yield, borrow, airdrop, burn, stake, liquidity, lend, loan, referral, mint, incentive
Codebase
Code
update contract, upgradable
(Erik Weitenberg, TNO; ope, 2023; Ethereum, 2023; oce, 2019; governance, 2023; Polygon, 2023)
Figure 6. Overlapping of Different Categories about Governance Issues.
@@ 3.2.2. Hierarchy of Governance Vulnerability
Figure 7. Hierarchy of Governance Issue.
@@ 3.2.3. Common Topics on Governance Issues
Table 5. Top-3 Topic Words in all governance issues.
@@
Gov Token
proposal, contract, token
voting power, power moved
governance, macilious, lack event
Ownership
centralization risk
trust issue admin, keys
blacklisted contract
Utility Token
token distribution initial
total supply, wrong total, supply restriction
rewards, price, users, duplicate, liquidity
Incentive
owner privileges, centralization risk
mint token
trust issue admin, keys
Revenue
protocol fee, transaction
interest rate, borrow, manipulate
incompatibility deationary tokens
Codebase
upgradeable contracts, storage
centralized control contract, contract upgrade
@@ 3.3.2. Detection of Inconsistency
Table 6. DeFi Tokenomics Configuration Inconsistency Checking, and F1 = 56.14%.